Calculate the freezing point of a solution made from 32.7 g of propane, C3H8, dissolved in 137.0 g of benzene, C6H6. The freezing point of benzene is 5.50°C and its Kf is 5.12°C/m.

1
Expert's answer
2022-04-15T03:22:03-0400

ΔTF = KF · b

ΔTF - the freezing-point depression, is defined as TF (pure solvent) − TF (solution)

KF - the cryoscopic constant

b - the molality


b = nsolute /  msolvent

b = n(C3H8) / m(C6H6) = (m(C3H8)/Mr(C3H8)) / m(C6H6) = (32.7 g / 44 g/mol) / 0.137 kg = 5.42 mol/kg


ΔT = 5.12°C/m * 5.42 mol/kg = 27.75°C

TF (solution) = ΔTF - TF (pure solvent)

Tf(C3H8) = ΔT - Tf(C6H6) = 27.75°C - 5.50°C = 22.25°C